Find the slope of the line passing through the pair of points (1/5,1) and 1/5,4)? My answer was undefined...
Found 2 solutions by RAY100, solver91311:
Answer by RAY100(1637)

Your answer was right.
.
a rough sketch show it best. A VERTICAL LINE at x= (1/5),,,,undefined
.
The formula also works, but is not a reassuring,,,m = (y2-y1) / (x2-x1) = 3/0 ,,,undefined

Answer by solver91311(24713)



Since the denominator of the slope fraction is the difference of the -coordinates, and the -coordinates of your given points are equal, the denominator of your slope fraction would be zero -- therefore undefined is exactly the correct answer.

John